# Orthogonal roots, quantum Hafnians, and generalized Rothe diagrams

### v1

@@ 4. Further examples of QP-triples
In this section, we give details of many other examples of QP-triples. Section 4.1 studies the root systems of rank $n$ where $n$-roots exist, and we show that for these root systems we can form a QP-triple by taking $U$ to be the entire set of positive roots (Theorem 4.3). In Section 4.2, we show that it is often possible to construct examples of QP-triples by selecting those positive roots in which particular simple roots appear with odd coefficients.
4.1. Orthogonal bases of roots
Recall from Lemma 2.3 that a Weyl group $W$ of rank $n$ and type $ADE$ has a set of $n$ orthogonal roots if and only if $W$ has type $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even, and we will show that $(W,S,{\Phi_{+}})$ is a QP-triple in these cases. These examples are particularly useful because many other examples of QP-triples can be constructed from them by restriction. The next result develops some key properties of residues of $n$-roots in these cases. Recall from Remark 3.5 (i) that for each $n$-root $R$ of $\Phi$, we have $\text{\rm Res}(R)=\text{\rm Res}_{\Phi_{+}}(R)$ and $\rho(R)=\rho_{\Phi_{+}}(R)$.
@@ 4.1. Orthogonal bases of roots
    The set of residues of $R$ can be expressed as the disjoint union
    $\text{\rm Res}(R)=\dot{\bigcup}_{Q\in D_{4}(R)}\text{\rm Res}(Q).$
@@
    $\text{\rm Res}(R)=\text{\rm Res}(Q)\quad\dot{\cup}\quad\dot{\bigcup}_{Q\subset H\in D_{6}(R)}(\text{\rm Res}(H)\backslash\text{\rm Res}(Q))\quad\dot{\cup}\quad\text{\rm Res}(R\backslash Q).$
@@
Remark 3.2 (i) and Proposition 2.9 (ii) imply that every $\gamma\in\text{\rm Res}(R)$ lies in the support of a unique $Q\in D_{4}(R)$. It follows that $\gamma\in\text{\rm Res}(Q)$ in this case, which proves $\text{\rm Res}(R)\subseteq\dot{\bigcup}_{Q\in D_{4}(R)}\text{\rm Res}(Q)$. Conversely, any element $\gamma\in\text{\rm Res}(Q)$ is fixed by all reflections $s_{\beta}$ such that $\Q\beta\in R\backslash$ Q, which implies that $\gamma\in\text{\rm Res}(R)$ and establishes the reverse containment. The disjointness assertion follows from Proposition 2.9 (iii), which completes the proof of (i).
@@
$\text{\rm Res}(H)=\text{\rm Res}(Q)\ \dot{\cup}\ \text{\rm Res}(Q^{\prime})\ \dot{\cup}\ \text{\rm Res}(Q^{\prime\prime})$
@@
$\text{\rm Res}(H)\backslash\text{\rm Res}(Q)=\text{\rm Res}(Q^{\prime})\ \dot{\cup}\ \text{\rm Res}(Q^{\prime\prime}).$
@@
By Lemma 2.3 and the paragraph preceding it, the set $\Omega_{\Phi_{+}}$ consists of the positive $n$-roots of $\Phi$, and the standard action of $W_{S}=W$ on $\Omega_{\Phi_{+}}$ is transitive. It is known [8, Theorem 4.5] that $\Omega_{\Phi_{+}}$ is a quasiparabolic set for $W$ under the standard action and level function $C+2N$, where $C$ and $N$ count the numbers of crossings and nestings in each $n$-root. Since $C(R)+2N(R)=\rho(R)=\rho_{\Phi_{+}}(R)$ for all $R\in\Omega_{\Phi_{+}}$ by Corollary 4.2, it follows that $(W,S,\Phi_{+})$ is a transitive QP-triple, which completes the proof. ∎
Corollary 4.4.
@@
This follows from Remark 3.12 and Proposition 4.7 of [8]. By that remark and Remark 3.5 (ii), if $\rho(s_{\alpha}Q)>\rho(Q)$ then we have (a) $Q$ is an alignment and $s_{\alpha}Q$ is a crossing, or (b) $Q$ is a crossing and $s_{\alpha}Q$ is a nesting, or (c) $Q$ is an alignment and $s_{\alpha}Q$ is a nesting. In these cases, the value of $\rho(s_{\alpha}(Q))-\rho(Q)$ is 1, 1, and 2, respectively, and $d$ is a nonnegative even number by parts (ii), (iii), and (iv) of [8, Proposition 4.7], respectively. ∎
Remark 4.5.
The paper [8] details many interesting properties of the quasiparabolic sets produced by Theorem 4.3. We highlight some of them that are directly related to this paper below.
  • (i)
  • (ii)
    The Poincaré polynomial of $PS_{\Phi_{+}}(q)$ factors into quantum integers, with the multiset of degrees being $\{3,5,\ldots,2k-1\}$ in type $D_{2k}$, $\{3,5,9\}$ in type $E_{7}$, and $\{3,5,9,15\}$ in type $E_{8}$ ([8, Proposition 7.1 (i)]). In particular, the cardinality of $\Omega_{\Phi_{+}}$ is $3\cdot 5\cdot 9=135$ in type $E_{7}$ and $3\cdot 5\cdot 9\cdot 15=2025$ in type $E_{8}$.
  • (iii)
    The unique $W$-minimal element in type $E_{8}$ is the set
    where $\theta_{4}$, $\theta_{6}$, $\theta_{7}$, and $\theta_{8}$ are the highest roots in the standard parabolic subsystems of types $D_{4}$, $D_{6}$, $E_{7}$, and $E_{8}$ of $\Phi$, respectively. This follows from [8, §6.3], and can also be checked directly using the definition of residues. Since the lowest-degree term of $PS_{\Phi_{+}}(q)$ is the constant 1 by (ii), the $8$-root $R$ has no residues with respect to $\Phi_{+}$.
@@ 4.2. An effective construction
The constructions of the following definition turn out to provide extremely effective methods for constructing QP-triples from the root systems of type $ADE$, and many of the resulting QP-triples provide constructions of known objects, in some cases providing extra insight into their structure. We will record the examples without giving any detailed proofs in this section, but we note that the examples in types $A$ and $D$ can all be interpreted in terms of the combinatorics of rook configurations, and we will study the examples $T(E_{8},8)$ and $T(E_{7},7)$ in full detail in Section 5.
@@
Example 4.7.
  • (i)
    The triple $(W,I,U)$ appearing in Theorem 3.10 is $T(D_{2k},2k)$. In type $D_{n}$, a simple root can only appear with coefficient $\pm 2$, $\pm 1$, or 0, so $U(D_{2k},2k)$ is the set of roots in which $\alpha_{2k}$ appears with coefficient $1$.
  • (ii)
    The triples $T(D_{2k},2k-1),T(D_{2k-1},2k-1),$ and $T(D_{2k-1},2k-2)\}$ are also QP-triples.
    The above three triples all give rise to essentially the same quasiparolic set as $T(D_{2k},2k)$, in the sense that their associated quasiparabolic sets $\Omega_{U(\Gamma,p)}$ all admit canonical level-preserving bijections to the quasiparabolic set $\Omega(D_{2k},2k)$ from (i). This bijection is induced by the symmetry of the Dynkin diagram for $\Omega(D_{2k},2k-1)$. For the case of $\Omega=\Omega(D_{2k-1},2k-1)$, each element of $\Omega$ corresponds to a (non-perfect) matching of $\{1,2,\dots,2k-1\}$ into (k1)(k-1) pairs and one unmatched point, and the canonical bijection from $\Omega$ to $\Omega(D_{2k},2k)$ sends such an element to the $2k$-root in $\Omega(D_{2k},2k)$ corresponding to the unique perfect matching containing those (k1)(k-1) pairs as blocks.
  • (iii)
    The triple $T(A_{n},p)$ is a QP-triple for all integers $n\geq 1$ and $1\leq p\leq n$.
    The residue sets arising from the triple $T(A_{n},p)$ resemble the Rothe diagrams mentioned in Remark 3.14 (i). More specifically, the set $U(A_{n},p)$ consists of the roots of the form $\varepsilon_{i}-\varepsilon_{j}$ where $1\leq i\leq p<j\leq n+1$, so we may identify $U$ with the positions on a $p\times(n+1-p)$ board with rows labeled by 1 to $p$ from top to bottom and columns labelled by $p+1$ to n+1pn+1-p from left to right. Each element of $R\in\Omega_{U(A_{n},p)}$ consists of $\min(p,n+1-p)$ orthogonal roots, and if we view $R$ as an involution in the natural way, as the product $w_{R}\in S_{n+1}$ of the 2-cycles $(i,j)$ for which $\varepsilon_{i}-\varepsilon_{j}\in R$, then a root $\varepsilon_{a}-\varepsilon_{b}\in U(A_{n},p)$ is a residue of $R$ with respect to $U(A_{n},p)$ if and only if we have both $w_{R}(a)<b$ and $w_{R}(b)>a$. Equivalently, a position $(a,b)$ on the board is a residue $R$ if and only if it is not attacked by a rook to its right or above it in the rook configuration corresponding to $R$, as shown in Figure 4, where the residues are marked by stars as in Figure 1.
@@
    Using the rook configuration model, it is possible to deduce that the Poincaré polynomial of $\Omega(A_{n},p)$ factors into quantum integers, as the product $PS_{U(A_{n},p)}=\prod_{D}[d]$ where $D$ is the set (with no repetition) containing each of the $\min(p,n+1-p)$ consecutive integers leading up to (and including) $\max(p,n+1-p)$. The row–column symmetry of the model also implies that there is a natural level-preserving bijection between the quasiparabolic sets $\Omega(A_{n},p)$ and $\Omega(A_{n},n+1-p)$; in particular, the Poincaré polynomials of $U(A_{n},p)$ and $U(A_{n},n+1-p)$ agree.
  • (iv)
    For all integers $n\geq 4$ and $1\leq p\leq n$, the triple $T(D_{n},p)$ is a QP-triple whenever we have $1\leq p\leq\lfloor n/2\rfloor$ (where $\lfloor\cdot\rfloor$ denotes the floor function), but not when $\lfloor n/2\rfloor<p<n-2$. However, if we relax the setting of Definition 4.6 and consider the set $U^{\prime}_{p}=U(D_{n},p)\cap U(D_{n},n)$, then the triple $T^{\prime}(\Gamma,P)=(W(D_{n}),S(D_{n})\setminus\{s_{p}\},U^{\prime}_{p})$ forms a QP-triple for all $1\leq p\leq n-2$. Note that the QP-triple $(W,I,U)$ from Theorem 3.13 is a special case of the triple $T^{\prime}(\Gamma,p)$ from (v), namely, $T^{\prime}(D_{2k},k)$.
    We note that for all the triples $T(D_{n},p)$ mentioned in the last paragraph, the elements $R\in\Omega_{U(D_{n},p)}$ and their residues can be interpreted in terms of rook configurations on a $p\times(n-p)$ board as in (iii), with the caveat that a position on the board may now correspond to either one of two roots $\varepsilon_{i}\pm\varepsilon_{j}$. The fact that $T(D_{n},p)$ fails to be a QP-triple when $\lfloor n/2\rfloor<p<n-2$ has to do with this caveat and the fact that in these cases $R$ does not contribute a rook to every row.
@@
  • (i)
    To summarize Example 4.7, every simple root $\alpha_{p}$ in a root system of type $A_{n}$ or $D_{n}$ gives rise to one or two QP-triples: $T(A_{n},p)$ in type $A_{n}$, $T^{\prime}(D_{n},p)$ in type $D_{n}$, and $T(D_{n},p)$ if $p\leq\lfloor n/2\rfloor$.
  • (ii)

### v2

@@ 4. Further examples of QP-triples
We give details of two families of examples of QP-triples in this section. The first family is associated to the root systems of rank $n$ where $n$-roots exist, and we show in Section 4.1 that for these root systems we can form a QP-triple by taking $U$ to be the entire set of positive roots.
In Section 4.2, we show that for each root system of simply-laced type $\Gamma$, it is often possible to construct a QP-triple $T(\Gamma,p)$ by taking $U$ to be the set of all positive roots in which a particular simple root $\alpha_{p}$ appears with odd coefficients (Definition 4.6). This construction furnishes a large number of QP-triples that form our second family of examples, and we will explain their rich connections to the combinatorics of perfect matchings, rook configurations, and Fano plane labellings. Two examples from this family, $T(E_{8},8)$ and $T(E_{7},7)$, also have natural connections to minuscule representations and del Pezzo surfaces, and a technical study of these two examples will be given in Section 5.
In Section 4.3, we explain how each QP-triple from sections 4.1 and 4.2 gives rise to a matroid, as well as how this matroid be used to predict the degrees of the QP-triple. We also explain how the examples $T(\Gamma,p)$ in the second family relate to symmetric spaces and symmetric pairs.
4.1. Orthogonal bases
Recall from Lemma 2.4 that a Weyl group $W$ of rank $n$ and type $ADE$ has a set of $n$ orthogonal roots if and only if $W$ has type $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even. We will show that $(W,S,{\Phi_{+}})$ is a QP-triple in these cases. These examples are particularly useful because many other examples of QP-triples can be constructed from them by restriction. The next result develops some key properties of the generalized Rothe diagrams of $n$-roots in these cases. Recall from Remark 3.6 (i) that for each $n$-root $R$ of $\Phi$, we have $\mathscr{D}(R)=\mathscr{D}_{\Phi_{+}}(R)$ and $\rho(R)=\rho_{\Phi_{+}}(R)$.
@@ 4.1. Orthogonal bases
    The generalized Rothe diagram $\mathscr{D}(R)$ can be expressed as the disjoint union
    $\mathscr{D}(R)=\dot{\bigcup}_{Q\in D_{4}(R)}\mathscr{D}(Q).$
@@
    $\mathscr{D}(R)=\mathscr{D}(Q)\quad\dot{\cup}\quad\left(\dot{\bigcup}_{Q\subset H\in D_{6}(R)}\mathscr{D}(H)\backslash\mathscr{D}(Q)\right)\quad\dot{\cup}\quad\mathscr{D}(R\backslash Q).$
@@
Remark 3.3 (i) and Proposition 2.10 (ii) imply that every $\gamma\in\mathscr{D}(R)$ lies in the support of a unique $Q\in D_{4}(R)$. It follows that $\gamma\in\mathscr{D}(Q)$ in this case, which proves $\mathscr{D}(R)\subseteq\dot{\bigcup}_{Q\in D_{4}(R)}\mathscr{D}(Q)$. Conversely, any element $\gamma\in\mathscr{D}(Q)$ is fixed by all reflections $s_{\beta}$ such that $\Q\beta\in R\backslash$ Q, which implies that $\gamma\in\mathscr{D}(R)$ and establishes the reverse containment. The disjointness assertion follows from Proposition 2.10 (iii), which completes the proof of (i).
@@
$\mathscr{D}(H)=\mathscr{D}(Q)\ \dot{\cup}\ \mathscr{D}(Q^{\prime})\ \dot{\cup}\ \mathscr{D}(Q^{\prime\prime})$
@@
$\mathscr{D}(H)\backslash\mathscr{D}(Q)=\mathscr{D}(Q^{\prime})\ \dot{\cup}\ \mathscr{D}(Q^{\prime\prime}).$
@@
By Lemma 2.4 and the paragraph preceding it, the set $\Omega_{\Phi_{+}}$ consists of the positive $n$-roots of $\Phi$, and the standard action of $W_{S}=W$ on $\Omega_{\Phi_{+}}$ is transitive. It is known [12, Theorem 4.5] that $\Omega_{\Phi_{+}}$ is a quasiparabolic set for $W$ under the standard action and level function $C+2N$, where $C$ and $N$ count the numbers of crossings and nestings in each $n$-root. Since $C(R)+2N(R)=\rho(R)=\rho_{\Phi_{+}}(R)$ for all $R\in\Omega_{\Phi_{+}}$ by Corollary 4.2, it follows that $(W,S,\Phi_{+})$ is a transitive QP-triple, which completes the proof. ∎
Remark 4.4.
  • (i)
    The Poincaré polynomial of $PS_{\Phi_{+}}(q)$ factors into quantum integers, with the multiset of degrees being $\{3,5,\ldots,2k-1\}$ in type $D_{2k}$, $\{3,5,9\}$ in type $E_{7}$, and $\{3,5,9,15\}$ in type $E_{8}$ ([12, Proposition 7.1 (i)]). We will explain how the degrees can be obtained combinatorially for type $D_{2k}$ in Example 4.7, and we will give a conceptual proof of the degrees in types $E_{7}$ and $E_{8}$, which were previously obtained only computationally, in the forthcoming paper [10]. In addition, we will also explain an algorithm for computing the degrees by using a matroid associated with $\Omega_{\Phi_{+}}$ in Section 4.3; see also Example 4.13.
  • (ii)
    Table 1 summarizes the key numerical quantities associated with the examples of Theorem 4.3, where $\kappa=n$ is the common cardinality of the sets in $\Omega_{U}$. The last column of the table concerns the matroid we can naturally associate to $\Omega_{\Phi_{+}}$, whose construction we will elaborate on in Section 4.3. Note that the cardinality of the set $\Omega_{U}$ is simply the product of the degrees by Definition 3.2 (iv), so $\lvert\Omega_{U}\rvert$ equals $(2k-1)!!,3\cdot 5\cdot 9=135$, and $3\cdot 5\cdot 9\cdot 15=2025$ in types $D_{2k},E_{7}$, and $E_{8}$, respectively.
    $\Gamma$ $\kappa$ $|U|$ degrees matroid
    $D_{2k}$ $2k$ 2k(2k1)2k(2k-1) $\{3,5,7,\ldots,2k-1\}$ $U^{(2)}_{k-1,k}$
    $E_{7}$ 77 $63$ $\{3,5,9\}$ PG(2,2)
    $E_{8}$ 88 $120$ $\{3,5,9,15\}$ AG(3,2)
    Table 1. QP-triples of the form $(W,S,{\Phi_{+}}_{+})$
  • (iii)
  • (iv)
    The unique $W$-minimal element in type $E_{8}$ is the set
    where $\theta_{4}$, $\theta_{6}$, $\theta_{7}$, and $\theta_{8}$ are the highest roots in the standard parabolic subsystems of types $D_{4}$, $D_{6}$, $E_{7}$, and $E_{8}$ of $\Phi$, respectively. This follows from [12, §6.3], and can also be checked directly using the definition of $\mathscr{D}(R)$. Since the lowest-degree term of $PS_{\Phi_{+}}(q)$ is the constant 1 by (ii), we have $\rho(R)=0$.
We close this subsection with the following lemma concerning the action of reflections on $\Omega_{\Phi_{+}}$. It will be used frequently in Section 5.
Lemma 4.5.
@@
This follows from Remark 3.12 and Proposition 4.7 of [12]. By that remark and Remark 3.6 (ii), if $\rho(s_{\alpha}Q)>\rho(Q)$ then we have (a) $Q$ is an alignment and $s_{\alpha}Q$ is a crossing, or (b) $Q$ is a crossing and $s_{\alpha}Q$ is a nesting, or (c) $Q$ is an alignment and $s_{\alpha}Q$ is a nesting. In these cases, the value of $\rho(s_{\alpha}(Q))-\rho(Q)$ is 1, 1, and 2, respectively, and $d$ is a nonnegative even number by parts (ii), (iii), and (iv) of [12, Proposition 4.7], respectively. ∎
@@ 4.2. An effective construction
The following definition produces a highly effective method for constructing QP-triples from simply-laced root systems. As our examples will show, many of the QP-triples constructed using the definition recover known objects, in some cases providing extra insight into their structure. The sets $U(\Gamma,p)$ in the definition correspond to symmetric spaces in a certain precise way that we will explain in Section 4.3.
@@
Note that when we expand a root in $\Phi$ into a linear combination of simple roots, the coefficient of each simple root lies in the set $\{\pm 1,0\}$ if $\Phi$ is of type $A$ and in the set $\{\pm 2,\pm 1,0\}$ if $\Phi$ has type $D_{n}$. Thus, in these two types, the condition in Definition 4.6 that $\alpha_{p}$ appears with odd positive coefficient is equivalent to the condition that $\alpha_{p}$ appears with coefficient 1.
In the rest of this subsection, we will describe all QP-triples of the form $T(\Gamma,p)$ where $\Gamma$ is of type $ADE$. We will explain how to interpret the examples in types $A$ and $D$ using the combinatorics of rook configurations, and we will study the examples $T(E_{8},8)$ and $T(E_{7},7)$ in full detail in Section 5. The key numerical quantities associated with the QP-triples are summarized in Table 2 below, where $U=U(\Gamma,p)$, $\kappa$ stands for the maximum cardinality of orthogonal subsets of $U$, $k$ stands for an integer larger than $1$, $n$ denotes the rank of $\Phi$, and $\lfloor\cdot\rfloor$ denotes the floor function. The Poincaré polynomial factors as a multiplicity-free product of quantum integers in all cases, so each example has a well-defined, multiplicity-free set of degrees; we will express this set in the notation
$a]_{b}:=\{a,a-1,a-2,...,a-b+1\}$
in some rows of the table. Recall that the cardinality of the quasiparabolic set $\Omega_{U}$ can be deduced from the degrees, as the product of the degrees, by Definition 3.2 (iv).
$\Gamma$ $p$ $\kappa$ $\lvert U\rvert$ degrees matroid
$A_{2k-1}$ $k$ $k$ $k^{2}$ $p]_{p-1}$ $U_{p-1,p}$
$A_{n}$ $n]\setminus\{(n+1)/2\}$ $m:=\min(p,n+1-p)$ m(n+1m)m(n+1-m) $n+1-m]_{m}$ $U_{m,m}$
$D_{2k}$ $k$ $2k$ $2k^{2}$ $p]_{p-1}$ $U^{(2)}_{p-1,p}$
$D_{n}$ $1,2,\ldots,\lfloor n/2\rfloor$ $2p$ 2p(np)2p(n-p) $n-p]_{p}$ $U^{(2)}_{p,p}$
$D_{2k}$ 2k1,2k2k-1,2k $k$ k(2k1)k(2k-1) $\{3,5,\ldots,2k-1\}$ $U_{k-1,k}$
$D_{2k+1}$ $2k,2k+1$ $k$ $k(2k+1)$ $\{3,5,\ldots,2k+1\}$ $U_{k,k}$
$E_{6}$ 22 44 2020 $\{2,3,5\}$ $AG(2,2)$
$E_{6}$ $1,6$ 22 $16$ $\{5,8\}$ $AG(1,2)$
$E_{7}$ $2,5$ 77 $35$ $\{2,3,5\}$ $PG(2,2)$
$E_{7}$ 11 44 $32$ $\{3,5,8\}$ $AG(2,2)$
$E_{7}$ 77 33 $27$ $\{5,9\}$ $PG(1,2)$
$E_{8}$ $1,2,5,6$ 88 $64$ $\{2,3,5,8\}$ $AG(3,2)$
$E_{8}$ 88 44 $56$ $\{5,9,14\}$ $AG(2,2)$
Table 2. QP-triples $T(\Gamma,p)$ of type ADEADE
Example 4.7 (Type $A$).
The triple $T(A_{n},p)$ is a QP-triple for all integers $n\geq 1$ and $1\leq p\leq n$. For such a triple, the set $U=U(A_{n},p)$ consists of the roots of the form $\varepsilon_{i}-\varepsilon_{j}$ where $1\leq i\leq p<j\leq n+1$, so we may identify $U$ with the positions on a $p\times(n+1-p)$ board with rows labeled by 1 to $p$ from top to bottom and columns labelled by $p+1$ to n+1pn+1-p from left to right. Each element of $R\in\Omega_{U(A_{n},p)}$ consists of $m:=\min(p,n+1-p)$ orthogonal roots, and if we view $R$ as an involution in the natural way, as the product $w_{R}\in S_{n+1}$ of the 2-cycles $(i,j)$ for which $\varepsilon_{i}-\varepsilon_{j}\in R$, then a root $\varepsilon_{a}-\varepsilon_{b}\in U(A_{n},p)$ dominates $R$ if and only if we have both $w_{R}(a)<b$ and $w_{R}(b)>a$. Equivalently, a position on the board lies in $\mathscr{D}_{U}(R)$ if and only if it is not attacked by a rook to its right or above it in the rook configuration corresponding to $R$, as shown in Figure 4, where the elements of $\mathscr{D}_{U}(R)$ are marked by stars as in Figure 1. Thus, the generalized Rothe diagrams associated to $T(A_{n},p)$ highly resemble the Rothe diagrams mentioned in Remark 3.15 (i).
@@
Using the rook configuration model, one can show using an elementary combinatorial argument that the Poincaré polynomial of $U(A_{n},p)$ is $\prod_{d\in[n+1-m]_{m}}[d]_{q}$, as asserted in Table 2. The row–column symmetry of the model implies that there is a natural level-preserving bijection between the quasiparabolic sets $\Omega(A_{n},p)$ and $\Omega(A_{n},n+1-p)$; in particular, the Poincaré polynomials of $U(A_{n},p)$ and $U(A_{n},n+1-p)$ agree.